OpenStudy (anonymous):

Valerie used to spend 10 h/wk watching TV. Now she spends less time per week watching TV. Which expressions represent how many hours per week Valerie spends watching TV now? Choose all answers that are correct. A. 10-(1/4 x 10) B. 10+(1/4*10) C. 3/4*10 D. 1/4*10

OpenStudy (retireed):

Your question is missing some important data. I think it is missing a 1/4 in the question.

OpenStudy (anonymous):

yea i forgot type it in

OpenStudy (retireed):

Based on that.... I see two correct answers. Remember it says to find Choose ALL answers that are correct. Which TWO answers equal the same number of hours.

OpenStudy (anonymous):

a and d

OpenStudy (retireed):

Fill in the blanks for ALL four answers. A. 10-(1/4 x 10) = ________ ?? B. 10+(1/4*10) = ________ ?? C. 3/4*10 = ________ ?? D. 1/4*10 = ________ ??

OpenStudy (anonymous):

look I cant do it I don't know it

OpenStudy (retireed):

Copy and paste my form into a post box Then overwrite the underlined section with the number of hours that the equation equals.

OpenStudy (anonymous):

I don't no how to do the questions

OpenStudy (anonymous):

@NetherCreep333

OpenStudy (retireed):

So basically you just guessed the answers earlier. Then you need to figure out how to calculate equations. Let's start one at a time. D is the easiest. D. 1/4*10 = ________ ?? since 1/4 equals 0.25. What is 0.25 times 10 ?

OpenStudy (anonymous):

2.5

OpenStudy (retireed):

Yes it is. Fill in the blanks for ALL four answers. A. 10-(1/4 x 10) = ________ ?? B. 10+(1/4*10) = ________ ?? C. 3/4*10 = ________ ?? D. 1/4*10 = 2.5 Now do C What is 3/4 ? Then take that number times 10.

OpenStudy (retireed):

Still waiting .... What is 3 divided by 4 ?

OpenStudy (retireed):

If 1/4 is 0.25 then what is 3 times 0,25 ?

OpenStudy (anonymous):

7.5

OpenStudy (anonymous):

brb

OpenStudy (retireed):

Okay Fill in the blanks for ALL four answers. A. 10-(1/4 x 10) = ________ ?? B. 10+(1/4*10) = ________ ?? C. 3/4*10 = 7.5 D. 1/4*10 = 2.5 Since you solved D, 1/4 * 10 is 2.5 You can solve A and B by adding 2.5 to 10 for answer A by subtracting 2.5 from 10 for answer B Once you have calculated the answers to A and B. See which TWO of the four answers are the same.

OpenStudy (anonymous):

back

OpenStudy (retireed):

OPPS i did that WRONG This correct Since you solved D, 1/4 * 10 is 2.5 You can solve A and B by subtracting 2.5 from 10 for answer A by adding 2.5 to 10 for answer B Once you have calculated the answers to A and B. See which TWO of the four answers are the same.

OpenStudy (anonymous):

a and d

OpenStudy (retireed):

A. 10-(1/4 x 10) = ________ ?? B. 10+(1/4*10) = ________ ?? This correct Since you solved D, 1/4 * 10 is 2.5 You can solve A and B by subtracting 2.5 from 10 for answer A by adding 2.5 to 10 for answer B

OpenStudy (retireed):

No that is the same guess you made before. SOLVE the equations, fill in the blanks. I have to go walk my dog. I will be back in an hour, so good luck.